# Medium effects on the electromagnetic form factors of $\rho$ meson

### v1

@@ VI.2 In-medium $\rho$ meson EMFFs
Now we turn to discuss the results of the $\rho^{+}$ meson EMFFs, $F_{1}^{\rho}(Q^{2})$, $F_{2}^{\rho}(Q^{2})$, and $F_{3}^{\rho}(Q^{2})$ in free space as well as in symmetric nuclear matter as depicted in Figs. 5(a)-(c). Results of $F_{1}^{*\rho}(Q^{2})$ for different nuclear matter densities as a function of $Q^{2}$ are illustrated in Fig. 5(a). At $\rho_{B}/\rho_{0}=0.0$ (free space), our result on $F_{1}^{\rho}(Q^{2})$ is consistent with that obtained in Ref. [54]. As the nuclear matter density and $Q^{2}$ increase, $F_{1}^{*\rho}(Q^{2})$ decreases. It is worth noting that $F_{1}^{*\rho}(Q^{2})$ always satisfies the normalization condition in free space and nuclear matter, i.e., $F_{1}^{\rho}(Q^{2}=0)=1.0$.
In Fig. 5(b), we show the results of the $F_{2}^{*\rho}(Q^{2})$ for different densities as a function of $Q^{2}$. We find that the results for $F_{2}^{*\rho}(Q^{2})$ decrease as the nuclear matter density and $Q^{2}$ increase. The decrease of $F_{2}^{*\rho}(Q^{2})$ is more pronounced at higher nuclear matter densities. We also find that in free space, our results are consistent with those reported in Ref. [54].
Results for the $F_{3}^{*\rho}(Q^{2})$ for different nuclear matter densities as a function of $Q^{2}$ are given in Fig. 5(c). Similar to the results of $F_{1}^{*\rho}(Q^{2})$ and $F_{2}^{*\rho}(Q^{2})$, we find that the $F_{3}^{*\rho}(Q^{2})$ decreases as the nuclear matter density $\rho_{B}/\rho_{0}$ and $Q^{2}$ increase. It is more significant at higher $\rho_{B}/\rho_{0}$. Again, our results of $F_{3}^{\rho}(Q^{2})$ in free space are consistent with those obtained in Ref. [54]. It is worth noting that in free space, $F_{3}^{\rho}(Q^{2}=0)$ equals the magnetic moment of the $\rho$ meson. The decrease of $F_{3}^{*\rho}(Q^{2}=0)$ at $\rho_{B}/\rho_{0}=1.0$ is approximately 0.15% in comparison to that for free space ($\rho_{B}/\rho_{0}=0.0$), which is insignificant reduction. At higher nuclear matter density, $\rho_{B}/\rho_{0}=2.0$, we find a reduction of the $F_{3}^{*\rho}(Q^{2}=0)$ by approximately 0.24% in comparison to that in free space, and again, insignificant.
@@
We are now in a position to present the in-medium $\rho^{+}$ meson EMFFs: the charge form factor $G_{C}^{\rho}(Q^{2})$ in units of the positron charge $e$, the magnetic form factor $G_{M}^{\rho}(Q^{2})$ in units of $e/(2m_{\rho})$, and the quadrupole form factor $G_{Q}^{\rho}(Q^{2})$ in units of $e/m_{\rho}^{2}$, both in free space and in nuclear matter. Results for the $\rho^{+}$ meson charge form factor for different nuclear matter densities as a function of $Q^{2}$ are illustrated in Fig. 6(a). We find that the values of $G_{C}^{\rho}(Q^{2})$ in free space are positive up to $Q^{2}\leq 3.2$ GeV2. At $Q^{2}\geq 3.4$ GeV2, the values of $G_{C}^{\rho}(Q^{2})$ become negative. At $Q^{2}=0$, $G_{C}^{\rho}(Q^{2})$ satisfies the charge normalization (charge conservation), namely, $G_{C}^{\rho}(0)=1.0$. This occurs not only in free space but also in a nuclear medium. At $\rho_{B}/\rho_{0}=1.0$, the positive values of $G_{C}^{*\rho}(Q^{2})$ slightly move up to $Q^{2}\leq 2.6$ GeV2, and for $Q^{2}\geq 2.8$ GeV2, the values of $G_{C}^{*\rho}(Q^{2})$ become negative, which starts smaller $Q^{2}$ value than that for free space. At higher nuclear matter densit $\rho_{B}/\rho_{0}=2.0$, the values of $G_{C}^{*\rho}(Q^{2})$ become positive for $Q^{2}\leq 2.2$ GeV2, while $G_{C}^{*\rho}(Q^{2})$ becomes negative at $Q^{2}\geq 2.4$ GeV2. Our result of $G_{C}^{\rho}(Q^{2})$ in free space is consistent with those reported in Ref. [54].
@@
Let us now discuss our results for the $G_{C}^{*\rho}(Q^{2})$ in the nuclear medium. Our results are consistent with those obtained in Ref. [52]. However, the differences are in the $Q^{2}$ values for crossing negative-positive values and the magnitudes. The results reported in Ref. [52] indicated that $G_{C}^{*\rho}(Q^{2})$ in nuclear matter is softer compared to our results. This is indicated by the onset of the negative values in $G_{C}^{*\rho}(Q^{2})$ occurring at a lower $Q^{2}$, i.e., at a smaller $Q^{2}$ value.
@@
Next, in Fig. 6(b), we show our predictions for $\rho^{+}$ meson magnetic moment form factors, $G_{M}^{*\rho}(Q^{2})$ for different values of $\rho_{B}/\rho_{0}$ and $Q^{2}$. We find that the $G_{M}^{*\rho}(Q^{2})$ decreases as the $\rho_{B}/\rho_{0}$ and $Q^{2}$ increase. This behavior of $G_{M}^{*\rho}(Q^{2})$ is consistent with the trend reported in Ref. [52], although the magnitudes differ. Note that the $G_{M}^{*\rho}(Q^{2})$ at $Q^{2}=0$ equals to the $\rho^{+}$ meson magnetic moment, $\mu_{\rho}^{*}$. In this work, we follow the formula of Ref. [13] in free space, where $\mu_{\rho}=G_{M}(0)\frac{M_{N}}{m_{\rho}}$. In the nuclear medium, the expression of $\mu_{\rho}$ is defined as $\mu^{*}_{\rho}=G_{M}^{*}(0)\frac{M_{N}^{*}}{m_{\rho}^{*}}$. The values of $m_{\rho}^{*}$ and $M_{N}^{*}$ for various nuclear matter densities are given in Table 1. We find that the value of $G_{M}^{\rho}(Q^{2})$ in free space at $Q^{2}=0$ is 2.082, in agreement with the value obtained in Ref. [54], while at $\rho_{B}/\rho_{0}=1.0$, we obtain $G_{M}^{*\rho}(0)=$ 2.079, which is very similar magnitude to that in free space. At higher density $\rho_{B}/\rho_{0}=2.0$, the value of $G_{M}^{*\rho}(0)$ is 2.077. The value of $G_{M}^{*\rho}(0)$ is 2.081 at $\rho_{B}/\rho_{0}=0.5$, which is smaller than the result obtained in Ref. [52], i.e., $G_{M}^{*\rho}(0)=2.18$ at $\rho_{B}/\rho_{0}=0.5$.
@@
Results for the quadrupole moment form factors of $\rho^{+}$ meson $G_{Q}^{*\rho}(Q^{2})$ for different nuclear matter densities as a function of $Q^{2}$ are illustrated in Fig. 6(c). We find that the $G_{Q}^{*\rho}(Q^{2})$ increases as the nuclear matter density and $Q^{2}$ increase. In free space, our result is consistent with that obtained in Ref. [13, 54]. Also, we obtain the same values of $G_{Q}^{\rho}(Q^{2})$ at $Q^{2}=0$ with those reported in Refs. [13, 54], i.e., $G_{Q}^{\rho}(0)=-0.87$ in free space. However, this value is rather different from that obtained in Ref. [52] in free space and also the value in nuclear matter. The difference is not only found in the values of $G_{Q}^{*\rho}(0)$, but also in the $Q^{2}$ dependence of $G_{Q}^{\rho}(Q^{2})$ for specific nuclear matter densities as shown in the lower-left panel of Fig. 5 in Ref. [52], where there is zero crossings in $G_{Q}^{*\rho}(Q^{2})$ for $\rho_{B}/\rho_{0}=0.00,0.25,$ and $0.50$ at low $Q^{2}$. Such behavior is not found in our results for the $G_{Q}^{*\rho}(Q^{2})$, as clearly shown in Fig. 6(c). We find that the values of the $G_{Q}^{*\rho}(0)$ for $\rho_{B}/\rho_{0}=0.5,1.0,1.5,$ and $2.0$ are 0.878,0.883,0.886,-0.878,-0.883,-0.886, and 0.888-0.888, respectively.
@@
In Fig. 6 (d), we show the results for the $\rho^{+}$ meson in-medium charge radius $\big{<}r_{\rho}^{*}\big{>}\equiv\big{<}r_{C}^{*2}\big{>}^{1/2}$ as a function of $\rho_{B}/\rho_{0}$. We find that the charge radius increases as the nuclear matter density increases. In free space, it is consistent with the value obtained in Ref. [54], i.e., $\big{<}r_{\rho}\big{>}=0.67$ fm. However, this is rather different from that obtained in Ref. [52], i.e., $\big{<}r_{\rho}\big{>}=0.52$ fm, as also clearly indicated in Fig. 6(d). Furthermore, our result in free space differs significantly from that of Ref. [13], where it was reported $\langle r_{\rho}\rangle=0.82$ fm. This discrepancy arises from the absence of pion loop contributions in the present work. Figure 6(d) also shows a comparison of our charge radius with that of Ref. [52]. The charge radius of Ref. [52] at $\rho_{B}/\rho_{0}=0.50$ (red filled asterisks) are close to our results (blue solid line). However, the results for $\big{<}r_{\rho}^{*}\big{>}$ are quite different from our results, as one can clearly see at $\rho_{B}/\rho_{0}=(0.0,$ and $0.90$). Note that, as already explained earlier, here we emphasize again that the $\rho_{B}/\rho_{0}=0.9$ corresponds to 0.84375 in the present study, as indicated in Fig. 6(d). For convenience, the values of the charge radius, quadrupole moment, and magnetic moment of the $\rho$ meson at various densities are listed in Table 1.
@@
Finally, we show the results of the $\rho^{+}$ meson magnetic and quadrupole moments for different nuclear matter densities, respectively, compared with those reported in Ref. [52], as depicted in Fig. 7(a)-(b). In Fig. 7(a), we show the result for the magnetic moment. We find that the $\rho^{+}$ meson magnetic moment decreases as the nuclear matter density increases. In addition, we find that our results for the magnetic moment are rather close to those obtained in Ref. [52]. The behavior of $\mu_{\rho}^{*}$ exhibits a similar trend to our prediction, as in Fig. 7(a), where the $\rho^{+}$ meson magnetic moment decreases with increasing nuclear matter density. We find $\mu_{\rho}^{*}=2.41\mu_{N}$ at $\rho_{B}/\rho_{0}=0.5$, whereas Ref. [52] reports $\mu_{\rho}^{*}=2.53\mu_{N}$ at the same density.
In Fig. 7(b), we show the results for the quadrupole moment of the $\rho^{+}$ meson as a function of the nuclear matter densities. We find that the $\mathcal{Q}_{\rho}^{*}$ increases as the nuclear matter density increases. Again, a similar trend is reported in Ref. [52], indicating the $\mathcal{Q}_{\rho}^{*}$ increases as the $\rho_{B}/\rho_{0}$ increases. At $\rho_{B}/\rho_{0}=0.0$ (free space) they obtained $\mathcal{Q}_{\rho}^{*}=-0.059$ fm2, which is close to our free space value, $\mathcal{Q}_{\rho}^{*}=-0.057$ fm2. However, at $\rho_{B}/\rho_{0}=0.5$ in their units with $\rho_{0}=0.15$ fm-3, namely 0.46785 in the present units, the values of $\mathcal{Q}_{\rho}^{*}$ is rather different. At this point, our calculation yields $\mathcal{Q}_{\rho}^{*}=-0.065$ fm2, whereas the value they obtained is $\mathcal{Q}_{\rho}^{*}=-0.088$ fm2, which is smaller than ours. As $\rho_{B}/\rho_{0}$ increases, the difference becomes more significant.

### v2

@@ VI.2 In-medium $\rho$ meson EMFFs
Now we turn to discuss the results of the $\rho^{+}$ meson EMFFs, $F_{1}^{\rho}(Q^{2})$, $F_{2}^{\rho}(Q^{2})$, and $F_{3}^{\rho}(Q^{2})$ in free space as well as in symmetric nuclear matter as depicted in Figs. 5(a)-(c). Results of $F_{1}^{*\rho}(Q^{2})$ for different nuclear matter densities as a function of $Q^{2}$ are illustrated in Fig. 5(a). At $\rho_{B}/\rho_{0}=0.0$ (free space), our result on $F_{1}^{\rho}(Q^{2})$ is consistent with that obtained in Ref. [Cloet:2014rja]. As the nuclear matter density and $Q^{2}$ increase, $F_{1}^{*\rho}(Q^{2})$ decreases. It is worth noting that $F_{1}^{*\rho}(Q^{2})$ always satisfies the normalization condition in free space and nuclear matter, i.e., $F_{1}^{\rho}(Q^{2}=0)=1.0$.
In Fig. 5(b), we show the results of the $F_{2}^{*\rho}(Q^{2})$ for different densities as a function of $Q^{2}$. We find that the results for $F_{2}^{*\rho}(Q^{2})$ decrease as the nuclear matter density and $Q^{2}$ increase. The decrease of $F_{2}^{*\rho}(Q^{2})$ is more pronounced at higher nuclear matter densities. We also find that in free space, our results are consistent with those reported in Ref. [Cloet:2014rja].
Results for the $F_{3}^{*\rho}(Q^{2})$ for different nuclear matter densities as a function of $Q^{2}$ are given in Fig. 5(c). Similar to the results of $F_{1}^{*\rho}(Q^{2})$ and $F_{2}^{*\rho}(Q^{2})$, we find that the $F_{3}^{*\rho}(Q^{2})$ decreases as the nuclear matter density $\rho_{B}/\rho_{0}$ and $Q^{2}$ increase. It is more significant at higher $\rho_{B}/\rho_{0}$. Again, our results of $F_{3}^{\rho}(Q^{2})$ in free space are consistent with those obtained in Ref. [Cloet:2014rja]. It is worth noting that in free space, $F_{3}^{\rho}(Q^{2}=0)$ equals the magnetic moment of the $\rho$ meson. The decrease of $F_{3}^{*\rho}(Q^{2}=0)$ at $\rho_{B}/\rho_{0}=1.0$ is approximately 0.15% in comparison to that for free space ($\rho_{B}/\rho_{0}=0.0$), which is insignificant reduction. At higher nuclear matter density, $\rho_{B}/\rho_{0}=2.0$, we find a reduction of the $F_{3}^{*\rho}(Q^{2}=0)$ by approximately 0.24% in comparison to that in free space, and again, insignificant.
@@
We are now in a position to present the in-medium $\rho^{+}$ meson EMFFs: the charge form factor $G_{C}^{\rho}(Q^{2})$ in units of the positron charge $e$, the magnetic form factor $G_{M}^{\rho}(Q^{2})$ in units of $e/(2m_{\rho})$, and the quadrupole form factor $G_{Q}^{\rho}(Q^{2})$ in units of $e/m_{\rho}^{2}$, both in free space and in nuclear matter. Results for the $\rho^{+}$ meson charge form factor for different nuclear matter densities as a function of $Q^{2}$ are illustrated in Fig. 6(a). We find that the values of $G_{C}^{\rho}(Q^{2})$ in free space are positive up to $Q^{2}\leq 3.2$ GeV2. At $Q^{2}\geq 3.4$ GeV2, the values of $G_{C}^{\rho}(Q^{2})$ become negative. At $Q^{2}=0$, $G_{C}^{\rho}(Q^{2})$ satisfies the charge normalization (charge conservation), namely, $G_{C}^{\rho}(0)=1.0$. This occurs not only in free space but also in a nuclear medium. At $\rho_{B}/\rho_{0}=1.0$, the positive values of $G_{C}^{*\rho}(Q^{2})$ slightly move up to $Q^{2}\leq 2.6$ GeV2, and for $Q^{2}\geq 2.8$ GeV2, the values of $G_{C}^{*\rho}(Q^{2})$ become negative, which starts smaller $Q^{2}$ value than that for free space. At higher nuclear matter density $\rho_{B}/\rho_{0}=2.0$, the values of $G_{C}^{*\rho}(Q^{2})$ become positive for $Q^{2}\leq 2.2$ GeV2, while $G_{C}^{*\rho}(Q^{2})$ becomes negative at $Q^{2}\geq 2.4$ GeV2. Our result of $G_{C}^{\rho}(Q^{2})$ in free space is consistent with those reported in Ref. [Cloet:2014rja].
@@
Let us now discuss our results for the $G_{C}^{*\rho}(Q^{2})$ in the nuclear medium. Our results are consistent with those obtained in Ref. [deMelo:2018hfw]. However, the differences are in the $Q^{2}$ values for crossing negative-positive values and the magnitudes. The results reported in Ref. [deMelo:2018hfw] indicated that $G_{C}^{*\rho}(Q^{2})$ in nuclear matter is softer compared to our results. This is indicated by the onset of the negative values in $G_{C}^{*\rho}(Q^{2})$ occurring at a lower $Q^{2}$, i.e., at a smaller $Q^{2}$ value.
@@
Next, in Fig. 6(b), we show our predictions for $\rho^{+}$ meson magnetic moment form factors, $G_{M}^{*\rho}(Q^{2})$ for different values of $\rho_{B}/\rho_{0}$ and $Q^{2}$. We find that the $G_{M}^{*\rho}(Q^{2})$ decreases as the $\rho_{B}/\rho_{0}$ and $Q^{2}$ increase. This behavior of $G_{M}^{*\rho}(Q^{2})$ is consistent with the trend reported in Ref. [deMelo:2018hfw], although the magnitudes differ. Note that the $G_{M}^{*\rho}(Q^{2})$ at $Q^{2}=0$ equals to the $\rho^{+}$ meson magnetic moment, $\mu_{\rho}^{*}$. In this work, we follow the formula of Ref. [Carrillo-Serrano:2015uca] in free space, where $\mu_{\rho}=G_{M}(0)\frac{M_{N}}{m_{\rho}}$. In the nuclear medium, the expression of $\mu_{\rho}$ is defined as $\mu^{*}_{\rho}=G_{M}^{*}(0)\frac{M_{N}^{*}}{m_{\rho}^{*}}$. The values of $m_{\rho}^{*}$ and $M_{N}^{*}$ for various nuclear matter densities are given in Table 1. We find that the value of $G_{M}^{\rho}(Q^{2})$ in free space at $Q^{2}=0$ is 2.082, in agreement with the value obtained in Ref. [Cloet:2014rja], while at $\rho_{B}/\rho_{0}=1.0$, we obtain $G_{M}^{*\rho}(0)=$ 2.079, which is very similar magnitude to that in free space. At higher density $\rho_{B}/\rho_{0}=2.0$, the value of $G_{M}^{*\rho}(0)$ is 2.077. The value of $G_{M}^{*\rho}(0)$ is 2.081 at $\rho_{B}/\rho_{0}=0.5$, which is smaller than the result obtained in Ref. [deMelo:2018hfw], i.e., $G_{M}^{*\rho}(0)=2.18$ at $\rho_{B}/\rho_{0}=0.5$.
@@
Results for the quadrupole moment form factors of $\rho^{+}$ meson $G_{Q}^{*\rho}(Q^{2})$ for different nuclear matter densities as a function of $Q^{2}$ are illustrated in Fig. 6(c). We find that the $G_{Q}^{*\rho}(Q^{2})$ increases as the nuclear matter density and $Q^{2}$ increase. In free space, our result is consistent with that obtained in Ref. [Carrillo-Serrano:2015uca, Cloet:2014rja]. Also, we obtain the same values of $G_{Q}^{\rho}(Q^{2})$ at $Q^{2}=0$ with those reported in Refs. [Carrillo-Serrano:2015uca, Cloet:2014rja], i.e., $G_{Q}^{\rho}(0)=-0.87$ in free space. However, this value is rather different from that obtained in Ref. [deMelo:2018hfw] in free space and also the value in nuclear matter. The difference is not only found in the values of $G_{Q}^{*\rho}(0)$, but also in the $Q^{2}$ dependence of $G_{Q}^{\rho}(Q^{2})$ for specific nuclear matter densities as shown in the lower-left panel of Fig. 5 in Ref. [deMelo:2018hfw], where there is zero crossings in $G_{Q}^{*\rho}(Q^{2})$ for $\rho_{B}/\rho_{0}=0.00,0.25,$ and $0.50$ at low $Q^{2}$. Such behavior is not found in our results for the $G_{Q}^{*\rho}(Q^{2})$, as clearly shown in Fig. 6(c). We find that the values of the $G_{Q}^{*\rho}(0)$ for $\rho_{B}/\rho_{0}=0.5,1.0,1.5,$ and $2.0$ are 0.878,0.883,0.886,-0.878,-0.883,-0.886, and 0.888-0.888, respectively.
@@
In Fig. 6 (d), we show the results for the $\rho^{+}$ meson in-medium charge radius $\big<r_{\rho}^{*}\big>\equiv\big<r_{C}^{*2}\big>^{1/2}$ as a function of $\rho_{B}/\rho_{0}$. We find that the charge radius increases as the nuclear matter density increases. In free space, it is consistent with the value obtained in Ref. [Cloet:2014rja], i.e., $\big<r_{\rho}\big>=0.67$ fm. However, this is rather different from that obtained in Ref. [deMelo:2018hfw], i.e., $\big<r_{\rho}\big>=0.52$ fm, as also clearly indicated in Fig. 6(d). Furthermore, our result in free space differs significantly from that of Ref. [Carrillo-Serrano:2015uca], where it was reported $\langle r_{\rho}\rangle=0.82$ fm. This discrepancy arises from the absence of pion loop contributions in the present work. Figure 6(d) also shows a comparison of our charge radius with that of Ref. [deMelo:2018hfw]. The charge radius of Ref. [deMelo:2018hfw] at $\rho_{B}/\rho_{0}=0.50$ (red filled asterisks) are close to our results (blue solid line). However, the results for $\big<r_{\rho}^{*}\big$> are quite different from our results, as one can clearly see at $\rho_{B}/\rho_{0}=(0.0,$ and $0.90$). Note that, as already explained earlier, here we emphasize again that the $\rho_{B}/\rho_{0}=0.9$ corresponds to 0.84375 in the present study, as indicated in Fig. 6(d). For convenience, the values of the charge radius, quadrupole moment, and magnetic moment of the $\rho$ meson at various densities are listed in Table 1.
@@
Finally, we show the results of the $\rho^{+}$ meson magnetic and quadrupole moments for different nuclear matter densities, respectively, compared with those reported in Ref. [deMelo:2018hfw], as depicted in Fig. 7(a)-(b). In Fig. 7(a), we show the result for the magnetic moment. We find that the $\rho^{+}$ meson magnetic moment decreases as the nuclear matter density increases. In addition, we find that our results for the magnetic moment are rather close to those obtained in Ref. [deMelo:2018hfw]. The behavior of $\mu_{\rho}^{*}$ exhibits a similar trend to our prediction, as in Fig. 7(a), where the $\rho^{+}$ meson magnetic moment decreases with increasing nuclear matter density. We find $\mu_{\rho}^{*}=2.41\mu_{N}$ at $\rho_{B}/\rho_{0}=0.5$, whereas Ref. [deMelo:2018hfw] reports $\mu_{\rho}^{*}=2.53\mu_{N}$ at the same density.
In Fig. 7(b), we show the results for the quadrupole moment of the $\rho^{+}$ meson as a function of the nuclear matter densities. We find that the $\mathcal{Q}_{\rho}^{*}$ increases as the nuclear matter density increases. Again, a similar trend is reported in Ref. [deMelo:2018hfw], indicating the $\mathcal{Q}_{\rho}^{*}$ increases as the $\rho_{B}/\rho_{0}$ increases. At $\rho_{B}/\rho_{0}=0.0$ (free space) they obtained $\mathcal{Q}_{\rho}^{*}=-0.059$ fm2, which is close to our free space value, $\mathcal{Q}_{\rho}^{*}=-0.057$ fm2. However, at $\rho_{B}/\rho_{0}=0.5$ in their units with $\rho_{0}=0.15$ fm-3, namely 0.46785 in the present units, the values of $\mathcal{Q}_{\rho}^{*}$ is rather different. At this point, our calculation yields $\mathcal{Q}_{\rho}^{*}=-0.065$ fm2, whereas the value they obtained is $\mathcal{Q}_{\rho}^{*}=-0.088$ fm2, which is smaller than ours. As $\rho_{B}/\rho_{0}$ increases, the difference becomes more significant.